# On the Fundamental Limitations of Dual Static CVaR Decompositions in Markov Decision Processes

### v1

@@ 3 An Explicit Mapping from the Value Function to the Static CVaR
While the static CVaR evaluation (Eq. 4) is inherently defined as a minimization problem, evaluating the value function of a risk-dependent policy (Eq. 6) requires $T$ recursive calls, hindering its interpretation as a well-defined minimization problem. In order to make explicit the optimization problem behind the latter, we first define the value function of a risk-dependent policy under a fixed set of state-level perturbations. For a fixed risk-dependent policy $\tilde{\pi}$, let $\bm{\tilde{\xi}}$ denote a complete specification of state-level perturbations such that $\bm{\tilde{\xi}}(\cdot|s,y,a)\in\Xi_{y}\left(P(\cdot|s,a)\right)$ for all states $s\in\mathcal{S}$, risk-levels $y\in[0,1]$, and actions $a=\tilde{\pi}(s,y)$. We further define $\bm{\tilde{\Xi}}(\tilde{\pi})$ as the set of all such valid state-level perturbations, highlighting the dependence on $\tilde{\pi}$.
Definition 5 (Policy-perturbation value function)
The value function of a risk-dependent policy $\tilde{\pi}$ under state-level perturbations $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}(\tilde{\pi})$ is the solution to:
@@
where we used action $a=\tilde{\pi}(s,y)$, updated risk-level $y^{\prime}=y\cdot\bm{\tilde{\xi}}(s^{\prime}|s,y,a)$, and $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}_{0}(s,y)=0$ for all states $s\in\mathcal{S}$ and risk-levels $y\in[0,1]$. We let $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s,y)\coloneq\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}_{T}(s,y)$.
This definition differs from the one in Equation 6 because it concerns fixed state-level perturbations $\bm{\tilde{\xi}}$ instead of computing them recursively at every iteration. The value function of a risk-dependent policy $\tilde{\pi}$ can now be seen as finding the best possible perturbation $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}(\tilde{\pi})$ to minimize this value.
@@
[Value function evaluation]lemmmalemmaDpOptimProblem The value function evaluation of a risk-dependent policy $\tilde{\pi}$ (Eq. 6) is equivalent to solving
$\bm{\upsilon}^{\tilde{\pi}}(s,y)=\min_{\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}\left(\tilde{\pi}\right)}\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s,y),$ (9)
where the above holds for all state-risk-level pairs $(s,y)$ simultaneously, meaning a single state-level perturbation policy $\bm{\tilde{\xi}}^{\star}$ is optimal for all $(s,y)$.
We now have two distinct optimization problems: the static evaluation over history perturbations $\xi$ (Eq. 4) and the value function evaluation over state-level perturbation policies $\bm{\tilde{\xi}}$ (Eq. 9). The two problems are in fact intimately connected. That is, any state-level perturbation $\bm{\tilde{\xi}}$ can be mapped to a corresponding history-level perturbation $\xi$ by taking the product of state-level perturbations along each history. For an initial risk-level $\alpha$ and history $H\in\mathcal{H}_{T}$, we define this mapping as
@@
where the risk levels $Y_{t}$ are incremented following $\bm{\tilde{\xi}}$ and starting from $Y_{0}=\alpha$. We now show that the mapping $\zeta_{\alpha}^{\bm{\tilde{\xi}}}$ is a valid history perturbation, and that it also recovers the value function for $\tilde{\pi}$ at the history-level.
@@
[State-level perturbation evaluation correspondence]propositionnpropositionConnexionDpCvar For any risk-dependent policy $\tilde{\pi}$, initial risk-level $\alpha$, and state-level perturbation $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}(\tilde{\pi})$, the mapped perturbation $\zeta_{\alpha}^{\bm{\tilde{\xi}}}$ is a valid history perturbation, that is $\zeta_{\alpha}^{\bm{\tilde{\xi}}}\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$), for which we have
@@
Because Proposition 5 applies to any state-level perturbation $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}(\tilde{\pi})$, in particular it applies to the optimal $\bm{\tilde{\xi}}^{\star}\in\arg\min_{\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}\left(\tilde{\pi}\right)}\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s_{0},\alpha)$. It directly follows that the value function evaluation (Lemma 9) is always an upper-bound to the static CVaR formulation over histories (Eq. 4).
Corollary 6 (Static CVaR upper-bound)
For any risk-dependent policy $\tilde{\pi}$ and initial risk-level $\alpha$, we have
@@
As a result of Corollary 6, the CVaR evaluation gap (Eq. 7) is non-zero if and only if the optimal history perturbation $\xi^{\star}$ is not in the image of the mapping $\zeta_{\alpha}$. That is, if the best global (history-level) perturbation cannot be decomposed into a sequence of valid local (state-level) perturbations, the value function evaluation will return an erroneous estimation of the policy’s static CVaR.
@@ 4 Characterizing the CVaR Evaluation Gap
We previously established that the value function of a risk-dependent policy provides an upper bound on its true static CVaR. In this section, we investigate the exact conditions under which the upper bound fails to be tight, leading to a CVaR evaluation gap. We show that the gap emerges when the optimal history-level perturbations are not realizable at the state level, a property that we formalize through a set of consistency constraints.
Definition 7 (Realizable trajectory perturbation)
For a given risk-dependent policy $\tilde{\pi}$ and initial risk-level $\alpha\in[0,1]$, a trajectory perturbation $\xi\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$ is realizable if there exists a state-level perturbation $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}(\tilde{\pi})$ such that $\zeta_{\alpha}^{\bm{\tilde{\xi}}}=\xi$.
The existence of such state-level perturbations $\bm{\tilde{\xi}}$ hinges on our ability to define a sequence of intermediate risk levels $Y_{t}$ that are mutually consistent for all histories. Recall that :-$\mathcal{H}\coloneq\bigcup_{t=1}^{T}\mathcal{H}_{t}$ is the set off all possible histories of length $|H|\leq T$ and define $H_{0:k}\coloneq(S_{0},A_{0},\dots,S_{k})\in\mathcal{H}_{k}$ as the $k$-length subsequence of a given history $H$. We formalize the mutual consistency notion by defining a risk-level assignment $\mathcal{Y}:\mathcal{H}\to[0,1]$ that maps any history $H$ to a risk-level $Y_{|H|}$. A risk-level assignment $\mathcal{Y}$ is consistent with respect to a trajectory perturbation $\xi$ if it enforces the correct total perturbation on all histories, while also respecting all stepwise constraints on the CVaR risk envelope and maintaining the correct sampled action from the risk-dependent policy. We will refer to these sets of constraints as the risk-assignment consistency constraints.
Definition 8 (Risk assignment consistency constraints)
For a given risk-dependent policy $\tilde{\pi}$, initial risk-level $\alpha\in[0,1]$, and history perturbations $\xi\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$, a risk-level assignment $\mathcal{Y}$ is consistent if it satisfies the following for all histories $H\in\mathcal{H}_{T}$ with $P^{\tilde{\pi}_{h}^{\alpha}}(H)>0$:
@@
[Consistency is necessary and sufficient for realizability]lemmmalemmaRealizabilityIffConsistency For any risk-dependent policy $\tilde{\pi}$ and initial risk-level $\alpha\in[0,1]$, history perturbations $\xi\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$ are realizable if and only if there exists a consistent risk-level assignment $\mathcal{Y}$ such that all risk-assignment consistency constraints (Def. 8) hold simultaneously.
The difficulty in satisfying the risk-assignment consistency constraints (Def. 8) lies in finding an assignment $\mathcal{Y}$ that satisfies all three constraint sets simultaneously. While it is clear that each constraint set can be satisfied in isolation, their intersection may be empty. This tension is the fundamental source of the CVaR evaluation gap (Eq. 7), which we formalize in the following theorem.
@@
[Conditions for CVaR evaluation gap]theoremthmAbsenceOfGap For any risk-dependent policy $\tilde{\pi}$ and initial risk-level $\alpha\in[0,1]$, we have $\operatorname{CVaR}_{\alpha}\left[Z^{\tilde{\pi}_{h}^{\alpha}}\right]=\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$ if and only if there exists at least one optimal history perturbations $\xi^{\star}$ solution to the static CVaR evaluation (Eq. 4) such that the risk-assignment constraints (Def. 8) have non-empty intersection.
Theorem 8 presents a formal characterization of necessary and sufficient conditions for when a CVaR evaluation gap occurs. It provides the valuable insight that, hidden under the mismatch between a risk-dependent policy’s value function and its static CVaR lies an unsolvable constraint satisfaction problem on the risk-level evolutions. More specifically, risk-dependent policies can induce action-selection consistency constraints that cannot be satisfied simultaneously with the other risk propagation and state-level risk envelope requirements, hampering the evaluation of a policy’s true CVaR.
The proposed constraint satisfaction perspective also clarifies why the evaluation is always accurate for risk-independent policies (Thm. 3.1 in Hau et al. (2023)). For such policies, the action-selection consistency constraint is non-binding, as the policy does not depend on the risk-level. A consistent risk-assignment can therefore always be constructed by recursively applying the CVaR decomposition Theorem (Thm. 2), guaranteeing the absence of a CVaR evaluation gap.
Corollary 9 (Existence of corresponding risk-dependent policy)
For any history-dependent policy $\pi_{h}:\mathcal{H}\to\mathcal{A}$, there exists a risk-dependent policy $\tilde{\pi}$ such that $\operatorname{CVaR}_{\alpha}\left[Z^{\pi_{h}}\right]=\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$ for all $\alpha\in[0,1]$.
The biggest issue that comes from Theorem 8 and Corollary 9 is that one cannot in general guarantee that all risk-dependent policies will have optimal history perturbations $\xi^{\star}$ that have consistent risk-assignment constraints (Def. 8). As a result, the set of all risk-dependent policies contains policies with a positive CVaR evaluation gap who will have an inaccurately high $\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$, impeding on the optimality of algorithms searching for $\tilde{\pi}^{\star}\in\arg\max_{\tilde{\pi}}\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$ like CVaR VI (Chow et al., 2015). The following section presents a worked example where the optimal risk-dependent policy has a positive CVaR evaluation gap and is therefore suboptimal.
A deeper dive into Hau et al. (2023)’s counterexample
@@
We now apply our newly introduced constraint satisfaction lens to a counterexample presented in Hau et al. (2023). We use the MDP shown in Figure 1, with horizon $T=2$ and initial risk $\alpha=0.5$. Note that our MDP differs superficially from the one in Hau et al. (2023), namely because we use a deterministic initial state $s_{0}$ with a single action available $a_{1}$ that does not impact the transition to the first state $S_{1}$, a procedure equivalent to the stochastic initial state presented in the original MDP.
For our example, we consider the risk-dependent policy $\tilde{\pi}$ produced by CVaR VI (Chow et al., 2015) and its corresponding optimal state-level perturbation $\bm{\tilde{\xi}}^{\star}$ solution to the risk-dependent-policy value function (Eq. 6):
@@
The optimal state-level perturbation $\bm{\tilde{\xi}}^{\star}$ therefore applies no changes to next states sampled, so the risk level remains $Y_{1}=0.5$ after the transition from $s_{0}$. As a result, the corresponding history-dependent takes action $a_{2}$ when reaching $s_{1}$, that is $\tilde{\pi}_{h}^{0.5}\left((s_{0},a_{1},s_{1})\right)=a_{2}$. Solving the static CVaR evaluation (Eq. 4), we find the corresponding history probabilities and optimal history perturbation $\xi^{\star}$:
$P^{\tilde{\pi}_{h}^{0.5}}(H)=\begin{cases}0.5,&H=(s_{0},a_{1},s_{1},a_{2},s_{5})\\ 0.5,&H=(s_{0},a_{1},s_{2},a_{1},s_{8})\end{cases}$ $\xi^{\star}(H)=\begin{cases}2,&H=(s_{0},a_{1},s_{1},a_{2},s_{5})\\ 0,&H=(s_{0},a_{1},s_{2},a_{1},s_{8})\end{cases}$
For the optimal history perturbation $\xi^{\star}$ to be realizable, there must exist a consistent risk-level assignment $\mathcal{Y}$ such that the risk-assignment consistency constraints (Def. 8) have non-empty intersection. Observing that histories of any length are fully defined by their final state since states never repeat in this MDP, the constraints can be expressed as:
@@
    Risk propagation: Directly applying trajectory perturbation $\xi^{\star}$, we get
@@
Figure 2 displays a visual representation of the risk-assignment constraints on the considered MDP (Figure 1) given the risk-dependent policy $\tilde{\pi}$ obtained using CVaR VI and an initial risk $\alpha=0.5$. We observe that the state-level risk envelope (Eq. 10) and the action-selection consistency (Eq. 11) constraints are impossible to satisfy simultaneously. Thus, no consistent risk-level assignment $\mathcal{Y}$ exists for $\xi^{\star}$. By Theorem 8, this confirms a positive CVaR evaluation gap at $\alpha=0.5$, explaining the empirical results obtained by Hau et al. (2023).

### v2

@@ 3 An Explicit Mapping from the Value Function to the Static CVaR
While the static CVaR evaluation (Eq. 5) is inherently defined as a minimization problem, evaluating the value function of a risk-dependent policy (Eq. 7) is defined as requiring $T$ recursive minimization problems, making for more cumbersome mathematical manipulations. In order to ease the manipulation of the latter, we first define the value function of a risk-dependent policy under a fixed set of state-level perturbations. For a fixed risk-dependent policy $\tilde{\pi}$, let $\bm{\tilde{\xi}}$ denote a complete specification of state-level perturbations such that $\bm{\tilde{\xi}}(\cdot|s,y,a)\in\Xi_{y}\left(P(\cdot|s,a)\right)$ for all states $s\in\mathcal{S}$, risk levels $y\in(0,1]$, and actions $a\in\mathcal{A}$. We further define $\bm{\tilde{\Xi}}\coloneq\left\{\bm{\tilde{\xi}}:\mathcal{S}\times(0,1]\times\mathcal{A}\to\mathcal{S}\times\mathbb{R}^{+}\,|\,\bm{\tilde{\xi}}(\cdot|s,y,a)\in\Xi_{y}\left(P(\cdot|s,a)\right)\forall(s,y,a)\in\mathcal{S}\times(0,1]\times\mathcal{A}\right\}$ as the set of all such valid state-level perturbations.
Definition 2 (Policy-perturbations value function).
The policy-perturbations value function of a risk-dependent policy $\tilde{\pi}$ under state-level perturbations $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}$ is obtained via:
@@
where we used action $a=\tilde{\pi}(s,y)$, updated risk level $y^{\prime}=y\cdot\bm{\tilde{\xi}}(s^{\prime}|s,y,a)$, and $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}_{0}(s,y)=0$ for all states $s\in\mathcal{S}$ and risk levels $y\in(0,1]$. We let $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s,y)\coloneq\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}_{T}(s,y)$.
This definition differs from the one in Equation 7 because it concerns fixed state-level perturbations $\bm{\tilde{\xi}}$ instead of computing them recursively at every iteration. The value function of a risk-dependent policy $\tilde{\pi}$ can now be seen as finding the best possible perturbations $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}$ to minimize this value.
@@
[Value function evaluation]lemmalemmaDpOptimProblem Under the conditions of Assumptions 1, the value function evaluation of a risk-dependent policy $\tilde{\pi}$ (Eq. 7) is equivalent to solving
$\bm{\upsilon}^{\tilde{\pi}}(s,y)=\min_{\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}}\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s,y),$ (10)
where the above holds for all state-risk level pairs $(s,y)$ simultaneously, meaning a single state-level perturbations set $\bm{\tilde{\xi}}^{\star}$ is optimal for all $(s,y)$.
Remark 3.
It is important to note that $\bm{\tilde{\Xi}}$ is the set of state-level perturbations $\bm{\tilde{\xi}}$ for the entire MDP, meaning that any difference in perturbation of a single (state, action, risk-level) tuple creates a different $\bm{\tilde{\xi}}$’. Therefore, taking the minimum over $\bm{\tilde{\Xi}}$ corresponds to selecting the worst possible simultaneous perturbation of the entire MDP for a given policy.
We now have two distinct single-step optimization problems for static CVaR evaluation: the static evaluation over history perturbations $\xi$ (Eq. 5) and the value function evaluation over state-level perturbations $\bm{\tilde{\xi}}$ (Eq. 10). The two problems are in fact intimately connected. That is, any state-level perturbations $\bm{\tilde{\xi}}$ can be mapped to corresponding history-level perturbations $\xi$ by taking the product of state-level perturbations along each history. For an initial risk level $\alpha$ and history $H\in\mathcal{H}_{T}$, we define this mapping as
@@
where the risk levels $Y_{t}$ are incremented following $\bm{\tilde{\xi}}$ and starting from $Y_{0}=\alpha$. We now show that the mapping $\zeta_{\alpha}^{\bm{\tilde{\xi}}}$ produces valid history perturbations recovering the value function for $\tilde{\pi}$ at the history-level.
@@
[State-level perturbations evaluation correspondence]propositionpropositionConnexionDpCvar Under the conditions of Assumptions 1, for any risk-dependent policy $\tilde{\pi}$, initial risk level $\alpha$, and state-level perturbations $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}$, the mapping $\zeta_{\alpha}^{\bm{\tilde{\xi}}}$ produces valid history perturbations, that is $\zeta_{\alpha}^{\bm{\tilde{\xi}}}\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$, for which we have
@@
Because Proposition 3 applies to any state-level perturbations set $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}$, in particular it applies to the optimal state-level perturbations set $\bm{\tilde{\xi}}^{\star}\in\arg\min_{\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}}\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s_{0},\alpha)$ for a fixed $\tilde{\pi}$. It follows that the value function evaluation (Eq. 10) is always an upper-bound to the true static CVaR of a policy (Eq. 5).
Corollary 1 (Static CVaR upper-bound).
Under the conditions of Assumptions 1, for any risk-dependent policy $\tilde{\pi}$ and initial risk level $\alpha$, we have
@@
As a result of Corollary 1, the CVaR evaluation gap (Eq. 8) is non-zero if and only if the optimal history perturbations $\xi^{\star}$ are not in the image of the mapping $\zeta_{\alpha}$. That is, if the best global (history-level) perturbations cannot be decomposed into a sequence of valid local (state-level) perturbations, the value function evaluation will return an erroneous estimation of the policy’s static CVaR.
@@ 4 Characterizing the CVaR Evaluation Gap
We established that the value function of a risk-dependent policy provides an upper bound on its true static CVaR. In this section, we now investigate the exact conditions under which the upper bound is strict, leading to a CVaR evaluation gap. We show that the gap emerges when the optimal history-level perturbations are not realizable at the state level, a property that we formalize through a set of consistency constraints.
Definition 3 (Realizable trajectory perturbations).
For a given risk-dependent policy $\tilde{\pi}$ and initial risk level $\alpha\in(0,1]$, trajectory perturbations $\xi\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$ are realizable if there exists state-level perturbations $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}$ such that $\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)=\xi(H)$ for all histories $H\in\mathcal{H}$.
The existence of such state-level perturbations $\bm{\tilde{\xi}}$ hinges on our ability to define a sequence of intermediate risk levels $Y_{t}$ that are mutually consistent for all histories. Recall that $\mathcal{H}\coloneq\bigcup_{t=1}^{T}\mathcal{H}_{t}$ is the set off all possible histories of length $|H|\leq T$ and define $H_{0:k}\coloneq(S_{0},A_{0},\dots,S_{k})\in\mathcal{H}_{k}$ as the $k$-length subsequence of a given history $H$. We formalize the mutual consistency notion by defining a risk level assignment $\mathcal{Y}:\mathcal{H}\to(0,1]$ that maps any history $H$ to a risk level $Y$. A risk level assignment $\mathcal{Y}$ is consistent with respect to trajectory perturbations $\xi$ if it enforces the correct total perturbations on all histories, while also respecting all stepwise constraints on the CVaR risk envelope and maintaining the correct sampled action from the risk-dependent policy. We will refer to these sets of constraints as the risk-assignment consistency constraints.
Definition 4 (Risk assignment consistency constraints).
For a given risk-dependent policy $\tilde{\pi}$, initial risk level $\alpha\in(0,1]$, and history perturbations $\xi\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$, a risk level assignment $\mathcal{Y}$ is consistent if, for all histories $H\in\mathcal{H}_{T}$ with $P^{\tilde{\pi}_{h}^{\alpha}}(H)>0$, it satisfies the following constraints:
@@
[Consistency if and only if realizability]lemmalemmaRealizabilityIffConsistency Under the conditions of Assumptions 1, , for any risk-dependent policy $\tilde{\pi}$ and initial risk level $\alpha\in(0,1]$, history perturbations $\xi\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$ are realizable if and only if there exists a consistent risk level assignment $\mathcal{Y}$ such that all risk-assignment consistency constraints (Def. 4) hold simultaneously.
The difficulty in satisfying the risk-assignment consistency constraints (Def. 4) lies in finding an assignment $\mathcal{Y}$ that satisfies all three constraint sets simultaneously. While it is clear that each constraint set can be satisfied in isolation, their intersection may be empty. This tension is the fundamental source of the CVaR evaluation gap (Eq. 8), which we formalize in the following theorem.
@@
[Conditions for CVaR evaluation gap]theoremthmAbsenceOfGap Under the conditions of Assumptions 1, for any risk-dependent policy $\tilde{\pi}$ and initial risk level $\alpha\in(0,1]$, we have $\operatorname{CVaR}_{\alpha}\left[Z^{\tilde{\pi}_{h}^{\alpha}}\right]=\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$ if and only if there exists at least one set of optimal history perturbations $\xi^{\star}$ solution to the static CVaR evaluation (Eq. 5) such that the risk-assignment constraints (Def. 4) can be satisfied simultaneously.
Theorem 4 presents a formal characterization of necessary and sufficient conditions for when a CVaR evaluation gap occurs. It provides the valuable insight that, hidden under the mismatch between a risk-dependent policy’s value function and its static CVaR lies an unsolvable constraint satisfaction problem on the risk level evolution. More specifically, risk-dependent policies can induce action-selection consistency constraints that cannot be satisfied simultaneously with the other risk propagation and state-level risk envelope requirements, hampering the evaluation of a policy’s true CVaR.
The proposed constraints satisfaction perspective also clarifies why the evaluation is always accurate for Markovian policies (Thm. 3.1 in Hau et al. (2023)). For such policies, the action-selection consistency constraints are non-binding, as the policy does not depend on the risk level. A consistent risk-assignment can therefore always be constructed by recursively applying the CVaR decomposition theorem (Thm. 1), guaranteeing the absence of a CVaR evaluation gap, as detailed in the following corollary.
{restatable}
[Existence of corresponding risk-dependent policy]corollarycorHistPolNoGap Under the conditions of Assumptions 1, for any Markovian policy $\pi:\mathcal{H}\to\mathcal{A}$ and initial risk level $\alpha\in(0,1]$, there exists a risk-dependent policy $\tilde{\pi}$ such that $\operatorname{CVaR}_{\alpha}\left[Z^{\pi}\right]=\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$.
The biggest issue that stems from Theorem 4 and Corollary 4 is that one cannot in general guarantee that all risk-dependent policies will have optimal history perturbations $\xi^{\star}$ that have consistent risk-assignment constraints (Def. 4). As a result, the set of all risk-dependent policies may contain policies with a positive CVaR evaluation gap who will have an inaccurately high $\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$, impeding on the optimality of algorithms searching for $\tilde{\pi}^{\star}\in\arg\max_{\tilde{\pi}}\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$ like CVaR VI (Chow et al., 2015). We now present a worked example where the optimal risk-dependent policy has a positive CVaR evaluation gap and is therefore suboptimal.
A deeper dive into the counterexample from Hau et al. (2023)
@@
We now apply our newly introduced constraint satisfaction analysis to a counterexample presented in Hau et al. (2023). We use the MDP shown in Figure 1, with horizon $T=2$ and initial risk $\alpha=0.5$. Note that our MDP differs superficially from the one in Hau et al. (2023), namely because we use a deterministic initial state $s_{0}$ with a single action available $a_{1}$ that does not impact the transition to the first state $S_{1}$, a procedure equivalent to the stochastic initial state presented in the original MDP.
For our example, we consider the risk-dependent policy $\tilde{\pi}$ produced by CVaR VI (Chow et al., 2015) and its optimal state-level perturbations $\bm{\tilde{\xi}}^{\star}$ solution to the risk-dependent-policy value function (Eq. 7):
@@
The optimal state-level perturbations $\bm{\tilde{\xi}}^{\star}$ therefore apply no changes to next states sampled, so the risk level remains $Y_{1}=0.5$ after the transition from $s_{0}$. As a result, the corresponding history-dependent policy takes action $a_{2}$ when reaching $s_{1}$, that is $\tilde{\pi}_{h}^{0.5}\left((s_{0},a_{1},s_{1})\right)=a_{2}$. Solving the static CVaR evaluation (Eq. 5), we find the corresponding history probabilities and optimal history perturbations $\xi^{\star}$:
$P^{\tilde{\pi}_{h}^{0.5}}(H)=\begin{cases}0.5&\text{if }H=(s_{0},a_{1},s_{1},a_{2},s_{5})\\ 0.5&\text{if }H=(s_{0},a_{1},s_{2},a_{1},s_{8})\end{cases}$ $\xi^{\star}(H)=\begin{cases}2&\text{if }H=(s_{0},a_{1},s_{1},a_{2},s_{5})\\ 0&\text{if }H=(s_{0},a_{1},s_{2},a_{1},s_{8})\end{cases}$
For the optimal history perturbations $\xi^{\star}$ to be realizable, there must exist a consistent risk level assignment $\mathcal{Y}$ such that the risk-assignment consistency constraints (Def. 4) have non-empty intersection. Observing that histories of any length are fully defined by their final state since states never repeat in this MDP, the constraints can be expressed as:
@@
    Risk propagation: Directly applying trajectory perturbations $\xi^{\star}$, we get
@@
Figure 2 displays a visual representation of the risk-assignment constraints on the considered MDP given the risk-dependent policy $\tilde{\pi}$ obtained using CVaR VI and initial risk level $\alpha=0.5$. We observe that the state-level risk envelope (Eq. 11) and the action-selection consistency (Eq. 12) constraints are impossible to satisfy simultaneously in state $s_{1}$. Thus, no consistent risk level assignment $\mathcal{Y}$ exists for $\xi^{\star}$. By Theorem 4, this confirms a positive CVaR evaluation gap, providing context to the results of Hau et al. (2023).
@@
Remark 4.
Not all risk-dependent policies have a CVaR evaluation gap. In Appendix C, we present such a risk-dependent policy on the same MDP from Figure 1 for which the dual-based DP evaluation accurately represents the static CVaR of its history-dependent counterpart.